Write a quadratic equation in the variable "x" having the given numbers as solutions.
Type the equation in standard form, ax^2 + bx + c = 0
(-sqrt( 7 ))(root ( 8, 7 ))

Answer by TakeATuition.com(57)

solution -
The form of quadratic equation is
ax^2 + bx + c = 0
if two roots of the equation are given then
equation can be formed by this formula
x^2 - (sum of the roots)x + product of the roots = 0 .....(i)

Now two roots are given that is 8 and 7
Therefore
sum of the roots = 8 + 7 = 15
product of the roots = 8*7 = 56
substituting the value in equation (i)
we have
x^2 - (15)x + 56 = 0
x^2 - 15x + 56 = 0
is the required equation.
FOR MORE HELP VISIT : www.TakeATuition.com